OpenStudy (anonymous):

Explain and show how you would estimate 0.5 x 4.62 . Then justify why this is a good choice for arriving at the correct answer.

OpenStudy (anonymous):

I would say 2.32 because when u multiply a number by 0.5, your answer is half the number :)

OpenStudy (anonymous):

2.31 actually

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its kay :)

OpenStudy (anonymous):

but how would i show it

OpenStudy (anonymous):

ermm divide each figure by 2?

OpenStudy (anonymous):

hmm mayby

OpenStudy (anonymous):

lol im not great at showing proof :(

OpenStudy (anonymous):

no i really bad at math so this is super hard for me i get numbers mixed up

OpenStudy (anonymous):

i'm

OpenStudy (e.cociuba):

0.5*4.62=4.62/2=2.31 It's as simple as that. Take 0.5 and multiply is by 4.62. Divide your answer by 2 and you get your answer.

OpenStudy (anonymous):

thank you mates

OpenStudy (e.cociuba):

Np!

OpenStudy (anonymous):

happy to help :)